Polarizers 1 and 3 have their axes of polarization indicated by the black solid lines, perpendicular to each other. Polarizer 2 has an axis of polarization that has an angle of e 70.2 with respect to the polarization axis of polarizer 1. What percentage of the initial (unpolarized) light intensity passes through? A: 1.66 B: 2.08 OC: 2.60 OD: 3.25 OE: 4.06 OF: 5.08 G: 6.35 H: 7.94. Submit Answer Tries 0/20show all work!!!
Answer #1

F) 5.08

Let Io is intensity of unpolaized light.

Intensity of light passing through first polarizer,

I1 = Io/2

Intensity of light passing through second polarizer,

I2 = I1*cos^2(70.2)

= Io*(1/2)*cos^2(70.2)

= Io*0.0574

Intensity of light passing through third polarizer,

I3 = I2*cos^2(90 - 70.2)

= Io*0.0574*cos^2(19.8)

= Io*0.0508

I3/Io = 0.0508

(I3/Io)*100 = 0.508*100

= 5.08%